OpenStudy (johnwillisneedshelp):

How many solutions does the equation −3y + 3y + 4 = 4 have?

jhonyy9 (jhonyy9):

so than you check this equation will result -3y+3y = 4-4 what mean 0 = 0 so indifferent what value you give to y the right side will be equal with left side so what mean this ? how many solutions has this equation ?

OpenStudy (mathmale):

I agree: If you simplify −3y + 3y + 4 = 4, you come up with 0 = 0. Think about what "solution" means here, before you try to respond to "how many solutions are there?"